OpenStudy (sissyedgar):

Another hard one! What is the solution to the equation? 3q + 5 = –2q + 20 A. q = 3 B. q = 5 C. q = 15 D. q = 25

OpenStudy (sissyedgar):

@paki

OpenStudy (sissyedgar):

uhh Paki

OpenStudy (sissyedgar):

u have A LOT of Medals... especially in History!!!!

OpenStudy (paki):

yeah.... i have..

OpenStudy (paki):

3q + 5 = –2q + 20 combine like terms....

OpenStudy (sissyedgar):

5?

OpenStudy (paki):

3q + 5 = –2q + 20 3q+2q = 20-5 solve now

OpenStudy (sissyedgar):

20-5=15...

OpenStudy (sissyedgar):

q = 15?

OpenStudy (paki):

yes 20-5 is 15 but 3q+2q = 5q so 5q 15... now solve for q

OpenStudy (sissyedgar):

yep, i just checked it! :P

OpenStudy (sissyedgar):

q = 15! Ty!

OpenStudy (paki):

so what u got

OpenStudy (sissyedgar):

^_^

OpenStudy (paki):

5q = 15 q = 3 .... check

OpenStudy (sissyedgar):

g2g